Population: {0,1,2}, n=2
Find the
A. mean of the population
B. the SD of the population
C. the sampling distribution of the sample mean
d. the mean of the probability distribution of means
e. the SD of the probability distribution of means
I have tried to work this out several times, but my answers don't match the answers in the back of the book. However, my teacher says that a lot of the answers in the back of the book are wrong, so I really have no idea if I am working these questions out right or not. Please show me how to work this problem so I can have a better understanding of what I am doing. Thanks

Answer by stanbon(57214)

Population: {0,1,2}, n=2
What do you mean by this?
It looks like you have a population of three numbers; but
then you say n=2 which usually means "the sample size is 2".
Is that the correct interpretation?
====================
Assuming it is I'll proceed:
A. mean of the population
(0+1+2)/3 = 3/3 = 1
------------------------
B. the SD of the population
0.8165
------------------------
C. the sampling distribution of the sample mean
For (0,1) x-bar = 0.5
For (0,2) x-bar = 1
for (1,2) x-bar = 3/2
-------------------------------
d. the mean of the probability distribution of means
????
e. the SD of the probability distribution of means
???
===========
Cheers,
Stan H.